Using elimination solve 4x-y=4 and 7x-3y=-3
Answer by stanbon(75887)

Using elimination solve 4x-y=4 and 7x-3y=-3
-------------------
4x-y = 4
7x-3y = -3
---------
28x - 7y = 7*4
28x - 12y = 4*-3
-------------------------
Subtract and solve for "y"::
5y = 40
y = 8
----
Solve for "x"::
4x-y = 4
4x -8 = 4
4x = 12
x = 3
----------
Cheers,
Stan H.
-----------